# AETAS: Analysis of Evolving Temporal Affect and Semantics for Legal History

### v1

@@ Abstract
Digital-humanities work on semantic shift often alternates between handcrafted close readings and opaque embedding machinery. We present a reproducible expert-system style pipeline that quantifies and visualises lexical drift in the Old Bailey Corpus (1720–1913), coupling interpretable trajectories with legally meaningful axes. We bin proceedings by decade with dynamic merging for low-resource slices, train skip-gram embeddings, align spaces through orthogonal Procrustes, and measure both geometric displacement and neighborhood turnover. Three visual analytics outputs, which are drift magnitudes, semantic trajectories, and movement along a mercy-versus-retribution axis, expose how justice, crime, poverty, and insanity evolve with penal reforms, transportation debates, and Victorian moral politics. The pipeline is implemented as auditable scripts so results can be reproduced in other historical corpora.
@@
Semantic drift , Diachronic embeddings , Digital humanities , Old Bailey , Expert systems[inst1]organization=PingCAP, Data & AI-Innovation Lab,city=Beijing,country=China
Highlights
Expert-system pipeline: decadal binning with adaptive merging, SGNS training, Procrustes alignment, and scripted visual analytics.
Digital-humanities storyline: connects lexical drift to penal transportation, the rise of professionalised justice, and Victorian moral care.
New interpretability device: a mercy-retribution axis quantifies value-laden shifts in legal vocabulary.
Reproducibility: all steps (binning, training, alignment, drift scoring, visualisation) are scripted for reruns on new corpora.
1 Introduction
Semantic drift studies in digital humanities now leverage distributional semantics to explain cultural change [1, 2]. Yet domain experts often face two gaps: (1) storylines that connect numeric drift to socio-legal events, and (2) tooling that behaves like an expert system with auditable steps and repeatable outputs. We address these gaps through AETAS (Analysis of Evolving Temporal Affect and Semantics; Latin aetas, meaning era or lifetime), a focused case study on the Old Bailey Corpus (OBC), a canonical source for London courtroom proceedings (1720–1913). The goal is to surface how legal and social concepts moved from early modern courts to late-Victorian professional justice.
We contribute: (i) a reproducible pipeline that treats diachronic embeddings as chained inference steps, (ii) a narrative that aligns semantic displacement with legal history, and (iii) visual artefacts tuned for interpretability rather than black-box accuracy.
@@ 2 Related work
Early semantic-change models used frequency- and context-based measures [1]. Subsequent surveys emphasise alignment stability, evaluation difficulties, and downstream interpretability [2, 3]. Contextualised models and probing have been explored for finer-grained senses [4]. Digital-humanities applications couple embeddings with historical argumentation, e.g., criminality, gender, and class [8, 11]. Our work extends this line with an expert-system stance: each step (binning, alignment, drift scoring) is explicit, audited, and paired with narrative-ready visualisations.
@@ 3.1 Corpus and temporal binning
We use OBC 2.0 TEI-XML, converting proceedings to a parquet corpus. Text is lowercased, punctuation stripped except hyphens, and grouped into 10-year bins. Bins with fewer than five million tokens merge forward (resulting in an initial 1674s bin through 1749), preserving chronological order.
@@ 3.2 Embedding training
Skip-gram with negative sampling (vector size 300, window 5, min count 10, negative samples 10, epochs 5, seed 42) is trained per bin using gensim [12]. Tokens per bin form sentence streams, avoiding shuffling across periods.
@@ 3.3 Space alignment
To compare meanings, we align each target space to a 1900s base using orthogonal Procrustes on the shared vocabulary [1]. Let and be base and target matrices; the rotation is obtained from with , ensuring distances are preserved.
@@ 3.4 Drift scoring and neighbors
Semantic drift between a word vector at start and at end is . We also track neighborhood turnover via Jaccard overlap between top- neighbors at the two times.
@@ 3.5 Normative axis
To ground interpretability, we build a mercy-retribution axis using averaged vectors of {mercy, pity, charity, kindness} minus {punishment, law, authority, order, duty}. Projections of target terms on this axis capture value-laden movement beyond raw drift.
@@ 3.6 Reproducibility
All stages (corpus parsing, binning, SGNS training, alignment, drift scoring, figure generation) are scripted for repeatable execution on other corpora; code and configuration are available in the project repository.
@@ 3.7 Corpus diagnostics
Table 1 reports decadal statistics. Early sparse decades (1674–1749) merge to satisfy the 5M-token rule; peak coverage is 1840s (13.2M tokens, 85K types). A naive lemmatisation heuristic reduces vocabulary size by 11–15% across decades, indicating moderate orthographic fragmentation without dominating the counts.
@@ 4.1 Drift magnitudes
Figure 1 shows cosine drift per target word across spans. On average, drift is , with the largest shifts in justice (0.57 in 1750s–1850s) and transportation (0.53), reflecting the decline of deportation and the rise of professional judges. Social terms such as charity exhibit long-horizon drift (0.50) as philanthropy professionalises.
@@ 4.2 Semantic trajectories
Figure 2 plots PCA projections of aligned vectors. Justice moves from a dispersed early-modern sense (neighbors include baron, recorder) toward a tight legal-professional cluster (bigham, phillimore). Poverty shifts away from moral failing toward structural conditions (neighbors like ill-health, schemes in 1900s).
@@ 4.3 Value-laden movement
Figure 3 shows projections on the mercy-retribution axis. Justice drifts slightly toward a more punitive neighbourhood by 1900s (score ), consistent with late-Victorian emphasis on deterrence despite procedural reforms. Insanity and poverty trend toward the merciful side by 1850s before softening in 1900s, echoing medicalisation and social-reform rhetoric. Figure 5 adds seed-variation bands; the sign of justice is robust to leave-one-out seeds, while poverty shows wider sensitivity because it co-occurs with both moralising and structural contexts.
@@ 4.4 Robustness checks
To mitigate brittleness, we enlarged neighborhoods to (Figure 4). Jaccard overlap rises modestly (mean 0.044 vs. 0 for ), indicating that low overlap is not solely a small- artifact. Target-word frequencies per million tokens stay within the same order across bins (e.g., justice 68290), suggesting drift is not driven by single bursts. A naive lemmatisation lowers type counts by 11–15%, implying that spelling variation inflates vocabulary moderately but does not dominate.
4.5 Contextual, CCA, VecMap, and pivot baselines
@@
We ran a contextual BERT-base baseline (up to 120 contexts/bin; early bins are sparse) with APD over contextual embeddings and substitute-based scaled JSD. Transportation shows the largest APD and JSD in 1750s–1850s, mirroring the SGNS signal; justice APD is moderate. A CCA alignment baseline (50 components) yields drift for justice and transportation that aligns with SGNS trends, with mean canonical correlations 0.63–0.68 over 6k–10k shared types. VecMap unsupervised alignment (capped vocabulary) produces consistent rankings (e.g., justice high drift in 1750s–1850s). To avoid explicit alignment, we compute a pivot baseline: rank similarity to 500 high-frequency pivots per bin and compare top-50 lists; pivot Jaccard remains low for justice but higher for crime/poverty, with mid-level Spearman stability. A temporal-reference norm baseline (anchor 1900s) also shows large norms for justice and transportation, consistent with long-range drift.
@@ 4.6 Incremental SGNS and multi-seed variance
We trained an incremental (time-forward) SGNS with three seeds over 1750s1850s1900s (vector size 200, 3 epochs per bin), saving snapshots per bin. Drift means (std across seeds) are: justice 0.40 (0.03) for 1750s–1850s, 0.50 (0.02) for 1750s–1900s; transportation 0.14 (0.01) for 1750s–1850s; crime 0.30 (0.01) for 1750s–1900s. Variance is low (std ) indicating seed robustness.

### v2

@@ Abstract
Digital-humanities work on semantic shift often alternates between handcrafted close readings and opaque embedding machinery. We present a reproducible expert-system style pipeline that quantifies lexical drift and its instability in the Old Bailey Corpus (1674–1913), coupling interpretable trajectories with legally meaningful axes. We bin proceedings by decade with dynamic merging for low-resource slices, train skip-gram embeddings, align spaces through orthogonal Procrustes to a 1900s anchor, and measure both geometric displacement and neighborhood turnover. We add split-half baselines and seed-sensitivity checks to separate within-bin instability from temporal change. Three visual analytics outputs (drift magnitudes, semantic trajectories, and movement along a mercy-versus-retribution axis) expose how justice, crime, poverty, and insanity evolve with penal reforms, transportation debates, and Victorian moral politics. The pipeline is implemented as auditable scripts so results can be reproduced in other historical corpora.
@@
semantic drift, diachronic embeddings, digital humanities, legal history, Old Bailey, interpretability, stability diagnostics1 Introduction
Semantic drift studies in digital humanities now leverage distributional semantics to explain cultural change (hamilton2016; kutuzov2018). Yet domain experts often face two gaps: (1) storylines that connect numeric drift to socio-legal events, and (2) tooling that behaves like an expert system with auditable steps and repeatable outputs. We address these gaps through AETAS (Analysis of Evolving Temporal Affect and Semantics; Latin aetas, meaning era or lifetime), a focused case study on the Old Bailey Corpus (OBC), a canonical source for London courtroom proceedings (1674–1913). The goal is to surface how legal and social concepts moved from early modern courts to late-Victorian professional justice.
A persistent obstacle is instability: embedding drift can reflect sampling variance or training noise as much as true semantic change. We therefore pair each drift claim with stability diagnostics (split-half baselines, seed sensitivity, and frequency controls), and report net drift/effect sizes that subtract within-bin instability before interpretation.
We contribute: (i) a reproducible pipeline that treats diachronic embeddings as chained inference steps, (ii) a stability-aware analysis that reports split-half baselines, seed sensitivity, anchor sensitivity, and net drift effect sizes, and (iii) a narrative with visual artefacts tuned for interpretability rather than black-box accuracy. Relative to prior SGNS+Procrustes pipelines, our emphasis is on stability diagnostics (aligned split-halves, permutation baselines), interpretability scaffolding (value axes with sensitivity and holdout checks), and explicit DH-ready exemplars. Compared to our earlier pipeline report, we add aligned split-half baselines with net-drift/effect sizes, anchor and seed sensitivity with chained alignment, hubness-aware overlap (CSLS) plus Global Anchors drift, permutation/downsampling controls, and axis sensitivity bands with a holdout sanity check.
@@ 2 Related work
Early semantic-change models used frequency- and context-based measures (hamilton2016). Subsequent surveys emphasise alignment stability, evaluation difficulties, and downstream interpretability (kutuzov2018; tahmasebi2021). Contextualised models and probing have been explored for finer-grained senses (giulianelli2020). Digital-humanities applications couple embeddings with historical argumentation, e.g., criminality, gender, and class (hill2019; underwood2019). Our work extends this line with an expert-system stance: each step (binning, alignment, drift scoring) is explicit, audited, and paired with narrative-ready visualisations. Dynamic and joint-training approaches reduce alignment variance by learning shared temporal spaces, including dynamic embeddings (bamler2017; rudolph2018). We retain independently trained SGNS models for auditability and interpretability, but compare against chain alignment and alignment-free baselines to contextualise these alternatives.
@@ 3.1 Corpus and temporal binning
We use OBC 2.0 TEI-XML, converting proceedings to a parquet corpus covering 1674–1913. OBC 2.0 is human-transcribed rather than OCRed, but titles, numerals, and proper-noun noise persist; we quantify their impact via NER/POS and proxy filters. Text is lowercased, punctuation stripped except hyphens, and grouped into 10-year bins. Bins with fewer than five million tokens merge forward (resulting in an initial 1674s bin through 1749), preserving chronological order.
@@ 3.2 Embedding training
Skip-gram with negative sampling (vector size 300, window 5, min count 10, negative samples 10, epochs 5, seed 42) is trained per bin using gensim (mikolov2013) with default subsampling (). Tokens per bin form sentence streams, avoiding shuffling across periods. We keep hyperparameters fixed across bins for comparability with standard diachronic SGNS settings; on 1,000-document samples, perturbations (window 2/8, dim 100, negative 5) preserve drift rankings (Spearman ), so we avoid per-bin tuning in sparse slices. A no-subsampling check on 2,000-document samples preserves drift rankings (Spearman across spans).
@@ 3.3 Space alignment
To compare meanings, we align each target space to a 1900s base using orthogonal Procrustes on the shared vocabulary (hamilton2016). Alignment uses the full shared vocabulary above the min-count threshold (10) in both bins; we do not cap the anchor set unless stated (e.g., Global Anchors). As a stable-anchor check, we select the 1,000 lowest-drift words among the top-20k shared types and re-align; drift rankings remain highly similar (Spearman , ). Because these anchors are identified after an initial alignment, we treat the stable-anchor variant strictly as a robustness probe rather than a primary result. Let and be base and target matrices; the rotation is obtained from with , ensuring distances are preserved. Mean-centering the matrices before Procrustes yields near-identical drift rankings (Spearman in across spans, –7), so we report uncentered results for consistency with prior work.
@@ 3.4 Drift scoring and neighbors
Semantic drift between a word vector at start and at end is . Neighbors are computed in the aligned spaces using the full vocabulary for each bin (no re-normalization is needed after orthogonal rotation). We track neighborhood turnover via Jaccard overlap between top- neighbors at the two times, .
3.5 Stability baselines and net drift
To quantify within-bin instability, we perform split-half training: for each bin, we randomly split texts into halves, train SGNS on each half (fixed seed to isolate sampling variance), align the two half spaces with orthogonal Procrustes, and measure within-bin drift for focal words. We repeat this 20 times and report mean and standard deviation; seed variance is checked separately via multi-seed incremental SGNS. For a cross-bin span , we define net drift as
where and are split-half means for the start and end bins. We also report a standardized effect size
using the split-half standard deviations where available. Net drift is interpreted as change that exceeds within-bin instability. We treat as a diagnostic rather than a significance test and additionally report seed-bootstrap uncertainty for drift (5 seeds on 2,000-document samples; median 95% CI width 0.054 across 19 word-span pairs). Table 3 lists net drift and values for all focal terms across the main spans.
@@ 3.6 Normative axis
To ground interpretability, we build a mercy-retribution axis using expanded seed sets of {mercy, pity, charity, kindness, lenity, forgiveness, clemency} minus explicitly punitive terms {punishment, penalty, execution, imprisonment, sentence, hanging, severity}, normalized to unit length. The axis is computed per time slice from that slice’s vectors. Projections of target terms on this axis (cosine to the axis) capture value-laden movement beyond raw drift. A holdout sanity check using non-seed mercy lexemes (e.g., pardon, grace, charitable, merciful) versus punitive lexemes (e.g., gaol, prison, conviction, discipline) is reported in the robustness section.
@@ 3.7 Reproducibility
All stages (corpus parsing, binning, SGNS training, alignment, drift scoring, figure generation) are scripted for repeatable execution on other corpora; the repository includes run-all.sh and a frozen environment spec (uv.lock, Python 3.12, gensim 4.3, transformers 4.57) with fixed random seeds. The repository URL and commit hash will be released at camera-ready; artifacts are available upon request. Because OBC is distributed under its own license, we do not redistribute raw TEI text; we will release derived drift tables, figures, and scripts to reconstruct results for licensed users. A minimal synthetic MRE (script + toy corpus) that runs end-to-end is included in the repository and can be shared during review upon request.
@@ 3.8 Corpus diagnostics
Table 1 reports decadal statistics. Early sparse decades (1674–1749) merge to satisfy the 5M-token rule; peak coverage is 1840s (13.2M tokens, 85K types). Headline drift spans begin at 1750s, with the early merged bin used only for chain-alignment diagnostics. A naive lemmatisation heuristic reduces vocabulary size by 11–15% across decades; a conservative orthography normalization on sampled bins reduces vocabulary by 6–7%, indicating moderate fragmentation without dominating the counts.
@@ 4.1 Drift magnitudes
Fig. 1 shows cosine drift per target word across spans. On average, drift is , with the largest shifts in justice (0.57 in 1750s–1850s) and transportation (0.53), reflecting the decline of deportation and the rise of professional judges. Social terms such as charity exhibit long-horizon drift (0.50) as philanthropy professionalises.
@@
To broaden the target set, we computed drift for the top 500 shared high-frequency words per span (frequency-controlled by minimum count across bins). The focal terms occupy high percentiles in this larger distribution: justice and transportation fall in the 99–100th percentile for 1750s–1850s, charity and justice are above the 98th percentile for 1750s–1900s, and most focal words sit above the 70th percentile for 1850s–1900s. This suggests the highlighted shifts are not artifacts of a tiny curated list, even though some targets (e.g., poverty in 1750s–1900s) are closer to the median. The full shared vocabularies contain 7,236 (1750s–1850s), 6,628 (1750s–1900s), and 10,446 (1850s–1900s) types above the min-count threshold; the repeated correlations correspond to seven focal words across three spans, with transportation falling below min-count in the 1900s (21 possible pairs 19 observed). Thus transportation is only reported for 1750s–1850s, and is excluded from 1750s–1900s/1850s–1900s tables and figures. To support independent scrutiny, we also export full ranked lists for the top-1,000 shared words per span, with frequency and sampled POS estimates. Windowed cue analysis around transportation indicates penal cues dominate while logistical cues are rare (penal-cue shares 0.07 in 1750s and 0.73 in 1850s; logistical cues ), and the term’s frequency collapses from 2,283 occurrences in 1750s to 62 in 1850s and 3 in 1900s, suggesting the drift reflects the disappearance of the penal sense rather than a shift to logistics.
@@ 4.2 Semantic trajectories
Fig. 2 plots PCA projections of aligned vectors. Justice moves from a dispersed early-modern sense (neighbors include baron, recorder) toward a tight legal-professional cluster (bigham, phillimore). Poverty shifts away from moral failing toward structural conditions (neighbors like ill-health, schemes in 1900s).
4.3 Close-reading exemplars
To ground the quantitative narrative, we sampled short excerpts from the corpus. In the 1750s bin, justice appears as an abstract goal: “…brought to justice, that I may be righted.” In the 1900s bin, it appears as a judicial title: “(BEFORE MR. JUSTICE BRAY.)” For poverty in the 1850s, mitigation language is explicit: “Recommended to mercy, in consequence of her poverty.” For insanity in the 1900s, the term is framed medically: “…in a state of distraction bordering on insanity, as the only way of ventilating his wrongs.” These snippets align with the measured shift toward professionalised court roles and medicalised explanations.
@@ 4.4 Value-laden movement
Fig. 3 shows projections on the mercy-retribution axis. Justice drifts slightly toward a more punitive neighbourhood by 1900s (score ), consistent with late-Victorian emphasis on deterrence despite procedural reforms. Insanity and poverty trend toward the merciful side by 1850s before softening in 1900s, echoing medicalisation and social-reform rhetoric. Fig. 6 adds seed-variation bands; the sign of justice is robust to leave-one-out seeds, while poverty shows wider sensitivity because it co-occurs with both moralising and structural contexts. Using a fixed axis defined in the 1900s and projecting earlier bins via alignment yields similar scores and a strong correlation with the per-slice axis (Spearman , ), suggesting the axis is not purely driven by seed drift. An antonym-pair axis (mercy–punishment, clemency–execution, lenity–severity, etc.) yields the same negative sign for justice in the 1900s (). A medicalisation–moralisation axis (medical vs. moral seeds) shifts insanity from negative in the 1750s () to positive by the 1850s/1900s (0.26), consistent with the narrative.
@@ 4.5 Robustness checks
To mitigate brittleness, we enlarged neighborhoods to (Fig. 4). Jaccard overlap rises modestly (mean 0.044 vs. 0.032 for ), indicating that low overlap is not solely a small- artifact. Using hubness-aware CSLS neighbors on a capped 10k vocabulary yields similarly low but slightly higher overlap (mean 0.028 for , 0.042 for ; Fig. 5). Rank-biased overlap (RBO, ) remains low (mean 0.047), and POS-restricted neighbors (NOUN/VERB/ADJ) increase slightly (mean 0.054), suggesting low overlap reflects genuine reconfiguration rather than only hubness or named entities. A qualitative audit of top neighbors clarifies the zero-Jaccard cases: in the 1750s, justice neighbors include magistrate, roundhouse, and deposition, whereas in the 1900s they are dominated by judge surnames such as bigham, grantham, and coleridge, indicating replacement rather than hubness artifacts. Removing judge surnames from neighbor lists leaves justice with zero Jaccard overlap across all spans, and other focal terms remain low (), so near-zero overlap is not driven solely by named-entity carryover. Anchor sensitivity tests show high agreement between alternative anchors and the 1900s anchor (Spearman for 1850s, for 1880s; 19 word-span pairs). A sequential chain alignment (1674s1910s) produces drift rankings consistent with the 1900s anchor (Spearman over 19 word-span pairs); starting the chain at 1750s (excluding the early mega-bin) yields essentially the same agreement (Spearman , ). Stable-anchor alignment (1,000 lowest-drift anchors among the top-20k shared types) yields similar rankings (Spearman , ). Downsampling 1850s/1900s bins to the 1750s token count yields moderate rank agreement with full-data drifts (Spearman for 1750s–1850s, for 1750s–1900s), suggesting that bin-size imbalance affects magnitude but not the highest-level ordering. Raising the merge threshold to 7M shifts early bins (start 1760s), but the 1850s–1900s drift ranking in sampled models matches the 5M baseline for overlapping terms (Spearman , ). On 1,000-document per-bin samples, hyperparameter perturbations (window 2/8, dim 100, negative 5) preserve rankings with Spearman in (). A coarse change-point scan over 1750s1810s1850s1900s shows the largest drift for justice, transportation, and charity in 1750s–1810s, while poverty and insanity peak in 1810s–1850s, aligning with early nineteenth-century reforms. A case-aware proper-noun proxy (dropping capitalized non-initial tokens, numerals, and honorific patterns) run on the main bins yields drift rankings highly correlated with the baseline (Spearman , ) and keeps justice/transportation among the highest drifts; mean overlap becomes 0.037 for and 0.040 for , so low overlap is not solely a named-entity artifact. A corpus-wide title filter shows justice tokens preceded by titles rise from 0.7% in the 1750s to 77–80% in the 1850s/1900s; training on the filtered corpus keeps drift large (0.51 for 1750s–1850s; 0.57 for 1750s–1900s) and retains a negative 1900s axis projection (), indicating the shift is not solely an honorific artifact. Using TEI judiciary-name lists to drop justice tokens followed by judge surnames removes 0.5%, 7.2%, and 46.7% of justice tokens in 1750s/1850s/1900s; drift remains high (0.61 for 1750s–1850s; 0.53 for 1750s–1900s), and the 1900s axis projection stays negative (). Seed-bootstrap CIs (5 seeds on 2,000-document samples) yield a median 95% CI width of 0.054 across 19 word-span pairs, indicating moderate uncertainty without overturning the main ordering. Time-shuffle permutation tests over all targets/spans yield values in , with justice above the null mean (1.32–1.53 SDs for 1750s–1850s/1900s) and transportation modestly above null (0.55 SDs in 1750s–1850s); we treat permutation results as a conservative sanity check rather than a significance claim. A frequency-matched random-seed null for the 1900s axis shows most target projections are not extreme (e.g., justice ), while charity is unusually high (). Target-word frequencies per million tokens stay within the same order across bins (e.g., justice 68290), suggesting drift is not driven by single bursts. Conservative lemmatisation/orthography normalization reduces type counts by 6–15%, implying that spelling variation inflates vocabulary moderately but does not dominate. A holdout axis sanity check finds non-seed mercy lexemes scoring higher than punitive lexemes in the 1900s bin (mean 0.07 vs. , vs. 12).
4.6 Contextual, CCA, VecMap, pivot, and Global Anchors baselines
@@
We ran a contextual BERT-base baseline (up to 120 contexts/bin; early bins are sparse) with APD over contextual embeddings and substitute-based scaled JSD. Transportation shows the largest APD and JSD in 1750s–1850s, mirroring the SGNS signal; justice APD is moderate. A CCA alignment baseline (50 components) yields drift for justice and transportation that aligns with SGNS trends, with mean canonical correlations 0.63–0.68 over 6k–10k shared types. VecMap unsupervised alignment (capped vocabulary) produces consistent rankings (e.g., justice high drift in 1750s–1850s). To avoid explicit alignment, we compute a pivot baseline: rank similarity to 500 high-frequency pivots per bin and compare top-50 lists; pivot Jaccard remains low for justice but higher for crime/poverty, with mid-level Spearman stability. A temporal-reference norm baseline (anchor 1900s) also shows large norms for justice and transportation, consistent with long-range drift. As an alignment-free global measure, Global Anchors (1,000 shared anchors) produces drift rankings correlated with SGNS (Spearman , ), again highlighting justice and transportation as high-shift terms. As an interpretable alternative, a dependency-slot JSD baseline on sampled contexts yields high divergence for justice and transportation in 1750s–1850s (JSD 0.65–0.69 where counts are sufficient), consistent with the embedding signal. Across spans, SGNS rank correlations with baselines are: Global Anchors in (–7), CCA in (–7), APD/JSD in where contextual counts permit (–5), and VecMap on the small span where it converged (). Pivot similarity is inversely related to drift (Spearman in ), as expected since lower Jaccard indicates greater change.
@@ 4.7 Incremental SGNS and multi-seed variance
We trained an incremental (time-forward) SGNS with three seeds over 1750s1850s1900s (vector size 200, 3 epochs per bin), saving snapshots per bin. Drift means (std across seeds) are: justice 0.40 (0.03) for 1750s–1850s, 0.50 (0.02) for 1750s–1900s; transportation 0.14 (0.01) for 1750s–1850s; crime 0.30 (0.01) for 1750s–1900s. Variance is low (std ) indicating seed robustness; this incremental model provides a simple dynamic baseline and preserves the main drift ordering for justice and transportation.